Entries and Schedules for Unfinished Jobs and Completed Jobs Tybee Industries Inc. uses a job order cost system. The followinMachine Hours 304 305 306 Total 227 f. Jobs completed: 301, 302, 303, and 305. g. Jobs were shipped and customers were billedMaterials ✓ Wages Payable Factory Overhead Accounts Payable d. Factory Overhead ✓ Accumulated Depreciation-Machinery and EquiCost of Goods Sold Finished Goods 2. Post the appropriate entries to T accounts for Work in Process and Finished Goods, usingNo. 306 Balance of Work in Process, January 30 4. Prepare a schedule of completed jobs on hand to support the balance in the
Answer #1

1)

Entries Description Debit credit
a Material 3090
Accounts payable 3090
b work in process 17120
Factory overhead 660
Material 17780
b-2 work in process 25820
Factory overhead 4450
wages payable 30270
c Factory overhead 3720
Accounts payable 3720
d Factory overhead 2140
Accumulated depreciation-machinery and equipment 2140
e work in process 10215
factory overhead(227*45) 10215
f Finished goods 29675
work in process 29675
g Accounts receivable 32380
sales 32380
cost of goods sold 19695
Finished goods 19695

2)

WORK IN PROCESS
B 42940     [17120+25820] F 29675 (balancing figure)
E 10215
Bal 23480
FINISHED GOODS
F 29675 G 19695 (balancing figure)
Bal 9980

3)

SCHEDULE OF UNFINISHED JOBS
Job no. Direct material Direct labor factory overhead Total
No.304 5340 7460 68*45=3060 15860
No.306 2480 3610 34*45= 1530 7620
Total 7820 11070 4590 23480

4)

SCHEDULE OF COMPLETED AND UNSOLD JOBS
Job no. Direct material Direct labor factory overhead Total
No.305 3390` 5690 20*45= 900 9980